OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. y=2/3x+9; (-6,5) A: y=-2/3x+1 B: y=-3/2x-4 C: y=-3/2x+3/2 D: y=2/3+9

OpenStudy (calculusxy):

perpendicular lines are always the negative (opposite) reciprocal of the current line's slope: \(\large \frac{2}{3}\) is now \(\large -\frac{3}{2}\) now substitute the x- and y-values and the slope of the perpendicular line into the slope-intercept formula: \(\large y= mx + b\) \(\large 5 = -\frac{3}{2}(-6) + b\) solve for \(\large b\)